# Fairness-Accuracy Trade-Offs: A Causal Perspective

### v1

@@ 1 Set $\lambda_{\text{low}}=0$, $\lambda_{\text{high}}=$ large
4       fit a neural network to solves the optimization problem in Eqs. 32-34 with $\lambda=\lambda_{\text{mid}}$ on $\mathcal{D}_{t}$ to obtain the predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$
@@
Definition 5 (Causal Lagrange Predictor $\widehat{Y}^{S}$).
The causally $S$-fair $\lambda$-optimal predictor $\widehat{Y}^{S}(\lambda)$ with respect to pathways in $S$ and the loss function $L$ is the solution to the following optimization problem:
$\operatorname*{arg\,min}_{f}$ $\;\mathbbm{E}\;L(Y,f(X,Z,W))+\lambda\big{(}\text{NDE}_{x_{0},x_{1}}(f)-\text{NDE}_{x_{0},x_{1}}(y)\cdot\mathbb{1}(\text{DE}\notin S)\big{)}^{2}+$ (32)
$\lambda\big{(}\text{NSE}_{x_{0}}(f)-\text{NSE}_{x_{0}}(y)\cdot\mathbb{1}(\text{SE}\notin S)\big{)}^{2}+\lambda\big{(}\text{NSE}_{x_{1}}(f)-\text{NSE}_{x_{1}}(y)\cdot\mathbb{1}(\text{SE}\notin S)\big{)}^{2}$ (34)
The above definition reformulates the problem of finding $\widehat{Y}^{S}$ to a Lagrangian form. This makes the problem amenable to standard gradient descent methods, and we propose a procedure for finding a suitable predictor $\widehat{Y}^{S}$ in Alg. 2 (CFCL). CFCL performs a binary search to find an appropriate value of the $\lambda$ parameter. For an interval $[\lambda_{\text{low}},\lambda_{\text{high}}]$ it takes the midpoint $\lambda_{\text{mid}}$. For this parameter value, it computes the optimal predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$ for the optimization problem in Eqs. 32-34 by fitting a feed-forward neural network with $n_{h}$ hidden layers and $n_{v}$ nodes in each layer. After this, the causal measures of fairness are computed for the predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$ on an evaluation set, and these measures are compared to the causal measures for the true outcome $Y$ (CFCL performs the appropriate hypothesis tests as in Eq. 31 ). If none of the hypotheses are rejected, it means that $\lambda_{\text{mid}}$ is large enough to enforce the causal fairness constraints, and the algorithm moves to the interval $[\lambda_{\text{low}},\lambda_{\text{mid}}]$. If a hypothesis is rejected, it indicates that $\lambda_{\text{mid}}$ is not large enough, and the algorithm moves to the interval $[\lambda_{\text{mid}},\lambda_{\text{high}}]$. Thus, CFCL leads to a systematic search for the parameter $\lambda$.
4 Experiments
In this section, we perform the causal fairness-accuracy analysis described in Sec. 2. We analyze the Census 2018 (Ex. 2), COMPAS (Ex. 3), and UCI Credit (Ex. 4 in Appendix E) datasets.
@@ 4 Experiments
Figure 4: Application of Alg. 1 on the Census 2018 dataset.
Example 2 (Salary Increase of Government Employees (Plečko and Bareinboim, 2024)).
The US government is building a tool for automated allocation of salaries for new employees. For developing the tool, they use the data collected by the United States Census Bureau in 2018, including demographic information $Z$ ($Z_{1}$ for age, $Z_{2}$ for race, $Z_{3}$ for nationality), gender $X$ ($x_{0}$ female, $x_{1}$ male), marital and family status $M$, education $L$, and work-related information $R$. The government wants to predict the outcome $Y$, the yearly salary of the employees (transformed to a log-scale). The standard fairness model (Fig. 3) is constructed as $\{X=X,Z=\{Z_{1},Z_{2},Z_{3}\},W=\{M,L,R\},Y=Y\}$.
@@
The analysis results are shown in Fig. 4, with uncertainty bars indicating standard deviations over 10 bootstrap repetitions. In the analysis of PSEL and TVD values (Fig. 4(a)), the team notices that imposing fairness constraints does not reduce RMSE, with the largest excess loss for the indirect effect, and smaller excess losses for direct and spurious effects. When looking at TVD values, they find that removing direct and indirect effects reduces the group differences substantially. In terms of causal fairness-utility ratios (Fig. 4(b)), the team finds that removing the direct effect has the best value in terms of reducing the disparity between groups vs. increasing the loss. The TV measure vs. excess loss dependence for different predictors $\widehat{Y}^{S}$ is shown in Fig. 4(c) (binary labels (D, I, S) in the figure indicate which effects were removed), and the graph $\mathcal{G}_{\text{PSEL}}$ with the values of PSEL and TVD for each transition is shown in Fig. 4(d). Based on the analysis, the team decides to use the predictor with the direct effect removed $\widehat{Y}^{D}$.
(a) (A)PSEL and (A)TVD values.
(b) CFUR values.
(c) Fairness/Utility of $\widehat{Y}^{S}$ predictors.
(d) $\mathcal{G}_{\text{PSEL}}$ with PSEL (blue) and TVD values (red).
Figure 5: Application of Alg. 1 on the COMPAS dataset.
Example 3 (Recidivism Prediction on COMPAS (Larson et al., 2016)).
The team from ProPublica wishes to understand the tension between fairness and accuracy, since the latter is an important concern of the district court. For the analysis, they use the complement of the area under the receiver operator characteristic curve ($1-$AUROC). They apply Alg. 1, and find that the APSEL value is the largest for the indirect effect (around 12% of AUROC), while it is smaller for the spurious effect (6%), and negligible for the direct effect, implying that the direct effect does not play a significant role in the prediction (see Fig. 5(a)). Regarding TVD values, the team finds that the removal of the spurious effect results in a 3% decrease in the TV measure, while the removal of the indirect effect decreases the disparity by 5%. Based on this, they find that the spurious effect is the best for reducing the disparity between groups (see Fig. 5(b) for CFUR values). Finally, the team also visualizes the TV and excess loss dependence (Fig. 5(c)), and the graph $\mathcal{G}_{\text{PSEL}}$ with PSEL and TVD values associated with each transition and effect removal (Fig. 5(a)). In the upcoming court hearing, the team proposes the usage of the causally-constrained predictor $\widehat{Y}^{\{D,S\}}$ that removes direct and spurious effects, and they use the above analysis to demonstrate the impact of this choice on the predictor’s accuracy.
@@ 5 Conclusion
The tension between fairness and accuracy is a fundamental concern in the modern applications of machine learning. The importance of this tension is also recognized in the legal frameworks of anti-discrimination, such as the disparate impact doctrine, which may allow for the usage of covariates correlated with the protected attribute if they are sufficiently important for the decision-maker’s utility (this concept is known as business necessity). In this work, we developed tools for analyzing the fairness-accuracy trade-off from a causal standpoint. Our approach allows the system designer to quantify how much excess loss is incurred when removing a path-specific causal effect (Def. 3). We also showed how the total excess loss, defined as the difference between the loss of the predictor fair along all causal pathways vs. an unconstrained predictor, can be decomposed into a sum of path-specific excess losses (Thm. 1). Based on this, we developed an algorithm for attributing excess loss to different causal pathways (Alg. 1), and introduced the notion of a causal fairness-utility ratio that captures the fairness gainexcess $\frac{\text{fairness gain}}{\text{excess loss}}$ ratio and in this way summarizes the trade-off for each causal path. Since our approach is dependent on causally-fair predictors (Def. 2), we introduced a new neural approach for constructing such predictors (Def. 5, Alg. 2). Finally, we demonstrated empirically that from causal perspective fairness and accuracy are almost always in tension (Exs. 2-4).
@@ Appendix A Linear Path-Specific Excess Losses
Note that we can compute:
$\mathbbm{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}$ $=\mathbbm{E}Y^{2}$ (39)
$=\mathbbm{E}[\alpha X+\gamma W+\epsilon_{y}]^{2}$ (40)
$=\mathbbm{E}[\epsilon_{y}^{2}]+\alpha^{2}\mathbbm{E}X^{2}+\gamma^{2}\mathbbm{E}W^{2}+2\mathbbm{E}\epsilon_{w}\mathbbm{E}\alpha X+\mathbbm{E}\epsilon_{w}\mathbbm{E}\gamma W$ (41)
$\quad+2\mathbbm{E}\alpha X\mathbbm{E}\gamma W$ (42)
$=\sigma_{y}^{2}+\frac{\alpha^{2}}{2}+\gamma^{2}\mathbbm{E}[\beta X+\epsilon_{w}]^{2}+2\frac{\alpha}{2}\mathbbm{E}\gamma[\beta X+\epsilon_{w}]$ (43)
@@
$\mathbbm{E}[Y-\widehat{Y}^{\text{DE}}]^{2}$ $=\mathbbm{E}[\alpha X+\gamma W+\epsilon_{y}-\gamma W]^{2}$ (46)
$=\sigma_{y}^{2}+\alpha^{2}\mathbbm{E}X^{2}$ (47)
@@
$\mathbbm{E}[Y-\widehat{Y}^{\text{IE}}]^{2}$ $=\mathbbm{E}[\alpha X+\gamma W+\epsilon_{y}-\alpha X]^{2}$ (49)
$=\sigma_{y}^{2}+\gamma^{2}\mathbbm{E}W^{2}$ (50)
@@
$\mathbbm{E}[Y-\widehat{Y}^{\emptyset}]^{2}$ $=\mathbbm{E}[\alpha X+\gamma W+\epsilon_{y}-\alpha X-\gamma W]^{2}$ (52)
$=\mathbbm{E}\epsilon_{y}^{2}=\sigma_{y}^{2}.$ (53)
Therefore, we have that
$\mathbbm{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}-\mathbbm{E}[Y-\widehat{Y}^{\emptyset}]^{2}$ $=\frac{\alpha^{2}+\gamma^{2}\beta^{2}+\alpha\gamma\beta}{2}+\gamma^{2}\sigma_{w}^{2}$ (54)
@@
$=\mathbbm{E}[Y-\widehat{Y}^{\text{DE}}]^{2}-\mathbbm{E}[Y-\widehat{Y}^{\emptyset}]^{2}$ (56)
$\quad+\mathbbm{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}-\mathbbm{E}[Y-\widehat{Y}^{\text{DE}}]^{2}.$ (57)
@@ Appendix B Theorem Proofs
$\text{PSEL}(\emptyset\to\{D,I,S\})=\mathbbm{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathbbm{E}[L(\widehat{Y}^{\emptyset},Y)].$ (58)
Let $S_{1},S_{2}\subset\{D,I,S\}$ such that $S_{1}\neq S_{2}$. The quantity can be expanded as follows:
$\text{PSEL}(\emptyset\to\{D,I,S\})$ $=\mathbbm{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathbbm{E}[L(\widehat{Y}^{\emptyset},Y)]$ (59)
$=\mathbbm{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathbbm{E}[L(\widehat{Y}^{S_{1}},Y)]$ (60)
$\quad+\mathbbm{E}[L(\widehat{Y}^{S_{1}},Y)]-\mathbbm{E}[L(\widehat{Y}^{\emptyset},Y)]$ (61)
$=\underbrace{\mathbbm{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathbbm{E}[L(\widehat{Y}^{\{S_{1},S_{2}\}},Y)]}_{\text{PSEL}(\{S_{1},S_{2}\}\to\{D,I,S\})}$ (62)
$\quad+\underbrace{\mathbbm{E}[L(\widehat{Y}^{\{S_{1},S_{2}\}},Y)]-\mathbbm{E}[L(\widehat{Y}^{S_{1}},Y)]}_{\text{PSEL}(S_{1}\to\{S_{1},S_{2}\})}$ (63)
$\quad+\underbrace{\mathbbm{E}[L(\widehat{Y}^{S_{1}},Y)]-\mathbbm{E}[L(\widehat{Y}^{\emptyset},Y)]}_{\text{PSEL}(\emptyset\to S_{1})}$ (64)
@@
Consider now the generalized $\mathcal{G}_{\text{PSEL}}$ graph for $m$ pathways shown in Fig. 6. The graph is split into columns, and in column $\ell$ all of the sets $S$ of size $\ell$ appear. The quantity $\text{APSEL}(S_{i})$ considers all pathw $\pi$ that start in $\emptyset$ and end in $\{S_{1},\dots,S_{m}\}$, and takes the average of excess losses associated with $S_{i}$ along each path $\pi$, written $\text{PSEL}(\pi^{<S_{i}}\to\pi^{\leq S_{i}})$. The APSEL quantity is written as:
@@
We now split the expression for the Shapley value by conditioning on the size of the set $S$:
$\phi(S_{i})$ $=\sum_{\ell=0}^{m-1}\sum_{S\subseteq\mathcal{S}\setminus\{S_{i}\}:|S|=\ell}\frac{1}{m\binom{m-1}{\ell}}\left(f(S\cup\{S_{i}\})-f(S)\right)$ (70)
$=\sum_{\ell=0}^{m-1}\sum_{S\subseteq\mathcal{S}\setminus\{S_{i}\}:|S|=\ell}\frac{1}{m\binom{m-1}{\ell}}\left(\text{PSEL}(\emptyset\to\{S\cup S_{i}\})-\text{PSEL}(\emptyset\to S)\right)$ (71)
@@
Now, fix a set $S$ of size $|S|=\ell$. There are exactly $\ell!$ paths leading from the $\emptyset$ to $S$. This is followed by a fixed transition $S\to S\cup S_{i}$ is chosen. The remaining steps from $S\cup S_{i}$ to $S_{1},\dots,S_{m}$ along the path $\pi$ can be chosen in $(m-\ell-1)!$ ways. Therefore, across all paths $\pi$, the contribution $\text{PSEL}(S\to\{S\cup S_{i}\}$ appears $\ell!(m-\ell-1)!$ times. Therefore, the APSEL quantity can be re-written as:
$\text{APSEL}(S_{i})$ $=\frac{1}{m!}\sum_{\ell=0}^{m-1}\sum_{S\subseteq\mathcal{S}\setminus\{S_{i}\}:|S|=\ell}{\ell!(n-\ell-1)!}\text{PSEL}(S\to\{S\cup S_{i}\})$ (73)

### v2

@@ 1 Set $\lambda_{\text{low}}=0$, $\lambda_{\text{high}}=$ large
4       fit a neural network to solves the optimization problem in Eqs. 52-56 with $\lambda=\lambda_{\text{mid}}$ on $\mathcal{D}_{t}$ to obtain the predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$
@@
The above definition reformulates the problem of finding $\widehat{Y}^{S}$ to a Lagrangian form. This makes the problem amenable to standard gradient descent methods, and we propose a procedure for finding a suitable predictor $\widehat{Y}^{S}$ in Alg. 2 (CFCL). In principle, many non-parametric learners such as boosting or neural networks could be used for fitting $\widehat{Y}^{S}$, and here we describe a neural approach. The key challenge for constructing $\widehat{Y}^{S}$ is finding the appropriate value of the tuning parameter $\lambda$. While the user may simply use a grid of values of $\lambda$, and inspect the loss function and the fairness measures, and then choose a $\lambda$ value, we propose a data-driven approach. We note that if $\lambda$ is too small, insufficient weight may be given to the fairness constraints, which therefore may be violated on new test data. If $\lambda$ is too high, however, we may give insufficient weight to minimizing the loss $L$, which may lead to poor performance on test data. Therefore, we propose a binary search type of procedure that first splits the data into train and evaluation folds, $\mathcal{D}_{t}$ and $\mathcal{D}_{e}$. CFCL starts with an interval $[\lambda_{\text{low}},\lambda_{\text{high}}]$ and takes the midpoint $\lambda_{\text{mid}}$. For this parameter value, it computes the optimal predictor $\widehat{Y}^{S}(\lambda_{\text{mid}})$ for the optimization problem in Eqs. 52-56 by fitting a feed-forward neural network with $n_{h}$ hidden layers and $n_{v}$ nodes in each layer. Then, for this fixed value of $\lambda_{\text{mid}}$, we test the hypotheses that
on evaluation data $\mathcal{D}_{e}$ (done in Eq. 57), which essentially test if the fairness constrains hold on the evaluation set $\mathcal{D}_{e}$, i.e., out of sample. If any of the hypotheses are rejected, it means that this value of $\lambda$ is too small to ensure that the fairness constraints are satisfied on new training data. Therefore, we want to find a larger $\lambda$, and the algorithm moves to the interval $[\lambda_{\text{mid}},\lambda_{\text{high}}]$. If none of the hypotheses are rejected, it means that $\lambda_{\text{mid}}$ is large enough to enforce the fairness constraints, and there may be an even smaller $\lambda$ that achieves this, so the algorithm moves to the interval $[\lambda_{\text{low}},\lambda_{\text{mid}}]$. In this way, CFCL leads to a data-driven way to select the tuning parameter $\lambda$. As the number of training and evaluation samples increases $|\mathcal{D}_{t}|,|\mathcal{D}_{e}|\to\infty$, the method is expected to perform increasingly well. An alternative approach would be to use a framework that automatically allows the learning of the $\lambda$ parameter (Fioretto et al. 2021).
@@
Figure 4: Application of Alg. 1 on the Census 2018 dataset. (a) Estimated APSEL (Eq. 40) and ATVD (Eq. 41) values; (b) The causal fairness-utility ratios (Eq. 51); (c) The Pareto plot for trade-offs between fairness (TV measure on vertical axis) and utility (excess RMSE on horizontal axis) for different predictors. The vector ($s_{1},s_{2},s_{3}$) indicates which of the DE/IE/SE pathways are constrained to zero; (d) The $\mathcal{G}_{\text{PSEL}}$ graph from Fig. 5 populated with PSEL and TVD values.
4 Experiments
In this section, we perform the causal fairness-accuracy analysis described in Sec. 2 on the Census 2018 dataset (Ex. 2). Additional analyses of the COMPAS (Ex. 3) and UCI Credit (Ex. 4) datasets are reported in Appendix E. All code for reproducing the experiments can be found in our Github repository https://github.com/dplecko/causal-acc-decomp.
Example 2 (Salary Increase of Government Employees (Plečko and Bareinboim 2024)).
The US government is building a tool for automated allocation of salaries for new employees. For developing the tool, they use the data collected by the United States Census Bureau in 2018, including
  • confounders $Z$, consisting of demographic information ($Z_{1}$ for age, $Z_{2}$ for race, $Z_{3}$ for nationality),
  • gender $X$ ($x_{0}$ female, $x_{1}$ male),
  • mediators $W$, including marital and family status $M$, education $L$, and work-related information $R$,
  • outcome $Y$, salary.
The government wants to predict the outcome $Y$, the yearly salary of the employees (transformed to a log-scale), in order to assign salaries for prospective employees. The standard fairness model (Fig. 2) is constructed as $\{X=X,Z=\{Z_{1},Z_{2},Z_{3}\},W=\{M,L,R\},Y=Y\}$.
@@ 4 Experiments
The analysis results are shown in Fig. 4, with uncertainty bars indicating standard deviations over 10 bootstrap repetitions. In the analysis of PSEL and TVD values (Fig. 4(a)), the team notices that imposing fairness constraints does not reduce RMSE for any of the effects. The largest excess loss is observed for the indirect effect, and smaller excess losses for direct and spurious effects. When looking at TVD values, they find that removing direct and indirect effects reduces the group differences substantially. In terms of causal fairness-utility ratios (Fig. 4(b)), the team finds that removing the direct effect has the best value in terms of reducing the disparity between groups vs. increasing the loss. The TV measure vs. excess loss dependence for different predictors $\widehat{Y}^{S}$ is shown in Fig. 4(c) (binary labels (D, I, S) in the figure indicate which effects were removed). The graph $\mathcal{G}_{\text{PSEL}}$ with the values of PSEL and TVD for each transition is shown in Fig. 4(d). Based on the analysis, the team realizes that it is possible to substantially reduce discrimination with a small amount of excess RMSE loss. They decide to implement the predictor $\widehat{Y}^{D}$ with the direct effect removed.
@@ 5 Conclusion
The tension between fairness and accuracy is a fundamental concern in the modern applications of machine learning. The importance of this tension is also recognized in the legal frameworks of anti-discrimination, such as the disparate impact doctrine, which may allow for the usage of covariates correlated with the protected attribute if they are sufficiently important for the decision-maker’s utility (this concept is known as business necessity). In this work, we developed tools for analyzing the fairness-accuracy trade-off from a causal standpoint. Our approach allows the system designer to quantify how much excess loss is incurred when removing a path-specific causal effect (Def. 3). We also showed how the total excess loss, defined as the difference between the loss of the predictor fair along all causal pathways vs. an unconstrained predictor, can be decomposed into a sum of path-specific excess losses (Thm. 1). Based on this, we developed an algorithm for attributing excess loss to different causal pathways (Alg. 1), and introduced the notion of a causal fairness-utility ratio that captures the
fairness gainexcess $\frac{\text{fairness gain}}{\text{excess loss}}$
ratio and in this way summarizes the trade-off for each causal path. Since our approach requires access to causally-fair predictors (Def. 2), we introduced a new neural approach for constructing such predictors (Def. 5, Alg. 2). Finally, we analyzed several real-world datasets, in order to investigate if fairness and utility are in a trade-off in practice. Our findings are that, from causal perspective, fairness and utility are almost always in tension (see Exs. 2-4), contrary to some other works appearing in the fair ML literature.
Acknowledgements
This research was supported in part by the NSF, ONR, AFOSR, DoE, Amazon, JP Morgan, and The Alfred P. Sloan Foundation.
@@ Appendix A Linear Path-Specific Excess Losses
Note that $\mathds{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}$ equals:
$=\mathds{E}Y^{2}$ (63)
$=\mathds{E}[\alpha X+\gamma W+\epsilon_{y}]^{2}$ (64)
$=\mathds{E}[\epsilon_{y}^{2}]+\alpha^{2}\mathds{E}X^{2}+\gamma^{2}\mathds{E}W^{2}+2\mathds{E}\epsilon_{w}\mathds{E}\alpha X$ (65)
$\quad+\mathds{E}\epsilon_{w}\mathds{E}\gamma W\quad+2\mathds{E}\alpha X\mathds{E}\gamma W$ (66)
$=\sigma_{y}^{2}+\frac{\alpha^{2}}{2}+\gamma^{2}\mathds{E}[\beta X+\epsilon_{w}]^{2}+2\frac{\alpha}{2}\mathds{E}\gamma[\beta X+\epsilon_{w}]$ (67)
@@
$\mathds{E}[Y-\widehat{Y}^{\text{DE}}]^{2}$ $=\mathds{E}[\alpha X+\gamma W+\epsilon_{y}-\gamma W]^{2}$ (70)
$=\sigma_{y}^{2}+\alpha^{2}\mathds{E}X^{2}$ (71)
@@
$\mathds{E}[Y-\widehat{Y}^{\text{IE}}]^{2}$ $=\mathds{E}[\alpha X+\gamma W+\epsilon_{y}-\alpha X]^{2}$ (73)
$=\sigma_{y}^{2}+\gamma^{2}\mathds{E}W^{2}$ (74)
@@
$\mathds{E}[Y-\widehat{Y}^{\emptyset}]^{2}$ $=\mathds{E}[\alpha X+\gamma W+\epsilon_{y}-\alpha X-\gamma W]^{2}$ (76)
$=\mathds{E}\epsilon_{y}^{2}=\sigma_{y}^{2}.$ (77)
Therefore, we have that $\mathds{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}-\mathds{E}[Y-\widehat{Y}^{\emptyset}]^{2}$ equals
$\frac{\alpha^{2}+\gamma^{2}\beta^{2}+\alpha\gamma\beta}{2}+\gamma^{2}\sigma_{w}^{2}$ (78)
@@
$=\mathds{E}[Y-\widehat{Y}^{\text{DE}}]^{2}-\mathds{E}[Y-\widehat{Y}^{\emptyset}]^{2}$ (80)
$\quad+\mathds{E}[Y-\widehat{Y}^{\{\text{DE, IE}\}}]^{2}-\mathds{E}[Y-\widehat{Y}^{\text{DE}}]^{2}.$ (81)
@@ Appendix B Theorem Proofs
$\text{PSEL}(\emptyset\to\{D,I,S\})=\mathds{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathds{E}[L(\widehat{Y}^{\emptyset},Y)].$ (82)
Let $S_{1},S_{2}\subset\{D,I,S\}$ such that $S_{1}\neq S_{2}$. Note that the quantity $\text{PSEL}(\emptyset\to\{D,I,S\})$ can be expanded as follows:
$\mathds{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathds{E}[L(\widehat{Y}^{\emptyset},Y)]$ (83)
$=\mathds{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathds{E}[L(\widehat{Y}^{S_{1}},Y)]$ (84)
$\quad+\mathds{E}[L(\widehat{Y}^{S_{1}},Y)]-\mathds{E}[L(\widehat{Y}^{\emptyset},Y)]$ (85)
$=\underbrace{\mathds{E}[L(\widehat{Y}^{\{D,I,S\}},Y)]-\mathds{E}[L(\widehat{Y}^{\{S_{1},S_{2}\}},Y)]}_{\text{PSEL}(\{S_{1},S_{2}\}\to\{D,I,S\})}$ (86)
$\quad+\underbrace{\mathds{E}[L(\widehat{Y}^{\{S_{1},S_{2}\}},Y)]-\mathds{E}[L(\widehat{Y}^{S_{1}},Y)]}_{\text{PSEL}(S_{1}\to\{S_{1},S_{2}\})}$ (87)
$\quad+\underbrace{\mathds{E}[L(\widehat{Y}^{S_{1}},Y)]-\mathds{E}[L(\widehat{Y}^{\emptyset},Y)]}_{\text{PSEL}(\emptyset\to S_{1})}$ (88)
@@
Consider now the generalized $\mathcal{G}_{\text{PSEL}}$ graph for $m$ pathways shown in Fig. 5. The graph is split into columns, and in column $\ell$ all of the sets $S$ of size $\ell$ appear. The quantity $\text{APSEL}(S_{i})$ considers all paths $\pi$ that start in $\emptyset$ and end in $\{S_{1},\dots,S_{m}\}$, and takes the average of excess losses associated with $S_{i}$ along each path $\pi$, written $\text{PSEL}(\pi^{<S_{i}}\to\pi^{\leq S_{i}})$. The APSEL quantity is written as:
@@
We now split the expression for the Shapley value $\phi(S_{i})$ by conditioning on the size of the set $S$:
$\sum_{\ell=0}^{m-1}\sum_{S\subseteq\mathcal{S}\setminus\{S_{i}\}:|S|=\ell}\frac{1}{m\binom{m-1}{\ell}}\left(f(S\cup\{S_{i}\})-f(S)\right)$ (94)
$=\sum_{\ell=0}^{m-1}\sum_{S\subseteq\mathcal{S}\setminus\{S_{i}\}:|S|=\ell}\frac{1}{m\binom{m-1}{\ell}}\big{(}\text{PSEL}(\emptyset\to\{S\cup S_{i}\})$ (95)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\text{PSEL}(\emptyset\to S)\big{)}$
@@
Now, fix a set $S$ of size $|S|=\ell$. There are exactly $\ell!$ paths leading from the $\emptyset$ to $S$. This is followed by a fixed transition $S\to S\cup S_{i}$ is chosen. The remaining steps from $S\cup S_{i}$ to $S_{1},\dots,S_{m}$ along the path $\pi$ can be chosen in $(m-\ell-1)!$ ways. Therefore, across all paths $\pi$, the contribution $\text{PSEL}(S\to\{S\cup S_{i}\}$ appears $\ell!(m-\ell-1)!$ times. Therefore, the APSEL$(S_{i})$ quantity can be re-written as:
$\frac{1}{m!}\sum_{\ell=0}^{m-1}\sum_{S\subseteq\mathcal{S}\setminus\{S_{i}\}:|S|=\ell}{\ell!(n-\ell-1)!}\text{PSEL}(S\to\{S\cup S_{i}\})$ (97)